A constant power P is acting on a car of mass m . The velocity increases from vand v2 .The distancce travelled by the car during this change is 
a,3p/m(v22-v12)
b,m/3p(v2-v1)
c,m/3p(v22-v12)
d,m/3p(v23-v13)

Grade:10

1 Answers

Vikas TU
7 years ago
Since Power is constant,
P = F x V => MaV 
and also,
a = vdv/dx
putting value of a in Power we get,
(P/m)dx = v^2dv
Integerating both sides we get,
Px/m = (v2^3 – v1^3)/3
x = (v2^3 – v1^3)m/3p